Discrete Structures problemSuppose we use p = 7 and q = 5 to generate keys for RSA. a) What is n ? b) What is on)? c) One choice of e is 5. What are the
Answer #1

a)

n= p*q

n=7*5 = 35

b)

Phi(n) = (p-1)(q-1)

Phi(n) = (7-1)(5-1)

Phi(n)=6*4 = 24

c) public key that is e should be coprime to phi(n) and should be less than phi(n).

Choice are 5,7,11,13,17,19,23

d) as explained in above part, (e,phi(n)) should be coprime and e<phi(n).

e) if e= 5, then to calculate d, we need to satisfy following equation.

=> (e*d) mod phi(n) = 1

=> ( 5*d) mod 24 = 1

Substitute d = 5, then above equation is satisfied.

So, d is 5

f) m = 3 (plain text)

To encrypt

m^e mod n

3^5 mod 35

243 mod 35 = 33

33 is cipher text.

g)

decryption

c^d mod n

33^5 mod 35 = 3

So, we got plain text back.